OpenStudy (anonymous):

The Ravens won twice as many games as they lost. They played 96 games. How many games did they win?

OpenStudy (ipwnbunnies):

Let 'x' be the amount of games they lost. Then, would you say '2x' is the games they won? Since they won twice as many as they lost?

OpenStudy (anonymous):

yes I think so

OpenStudy (ipwnbunnies):

Good. And 96 is the total number of games the Ravens played. That include their wins and losses, of course. We can create an equation. x + 2x = 96 Losses + Wins = total amount of games

OpenStudy (ipwnbunnies):

Solve for x. Plug 'x' back into the equation for the amount of games the Ravens won.

OpenStudy (anonymous):

okay then what? you just divide?

OpenStudy (ipwnbunnies):

Can you solve for x?

OpenStudy (anonymous):

divide 96 by 2? is that what I do?

OpenStudy (ipwnbunnies):

No. x + 2x = 96 You must combine similar terms first.

OpenStudy (anonymous):

oh yeah so 3x = 96 right?

OpenStudy (ipwnbunnies):

Right. Solve for 'x'. Plug it back into the equation we made for the number of games they won.

OpenStudy (anonymous):

what do you mean? wouldn't I just divide 3 by 96??

OpenStudy (ipwnbunnies):

Yes, to solve for x.

OpenStudy (anonymous):

okay so 96/3 = 32

OpenStudy (ipwnbunnies):

Right. x = 32 Earlier, we said 'x' is the number of games the Ravens lost. '2x' must be the number of games they won.

OpenStudy (anonymous):

which means what?

OpenStudy (anonymous):

that means they won 32 games?

OpenStudy (ipwnbunnies):

Look at your question, m8. They want you to find the number of games they won.

OpenStudy (anonymous):

yes they do

OpenStudy (ipwnbunnies):

I just said 'x' is the number of games they lost. .-.

OpenStudy (ipwnbunnies):

2x is the number of games they won.

OpenStudy (anonymous):

wait so that means 2(32) = 64 games won?

OpenStudy (ipwnbunnies):

YES. YESSSS. Gewd job. \[\heartsuit\]

OpenStudy (anonymous):

okay sorry im a little slow! :/

OpenStudy (anonymous):

Thanks for the help!